OpenStudy (anonymous):

What is the solution set of the absolute value sentence |3k + 4| = 8? A. {4/3, -4} B. {x|x > 4 or x < -4/5} C. No Solution D. {x|-11 < x < -7} E. All Real Numbers

OpenStudy (anonymous):

Set 3k+4 = 8 and -8 and solve each for x. :)

OpenStudy (anonymous):

|3k + 4| = 8 3k + 4 = 8 or 3k + 4 = - 8 3k = 4 k = 4/3 3k + 4 = - 8 3k = - 12 k = - 3

OpenStudy (anonymous):

Sorry, -4 in the last one.

OpenStudy (anonymous):

Was just about to mention that :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

You got it, @lisa?

OpenStudy (anonymous):

:D yes..... i'm working it out by my self just in case ana2 is wrong lol j/k :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Smart move. And that's not to be a crack on @ana :) ALWAYS check someone's work to 1) make sure you understand what they did, and 2) to make sure they didn't make mistakes. :) We all mistype at some point in time...

OpenStudy (anonymous):

lol true :)

OpenStudy (anonymous):

were all im perfect here lol

OpenStudy (anonymous):

imperfect

OpenStudy (anonymous):

lol